OpenStudy (anonymous):

Find the absolute maximum value for f(x)=x-(2/x^2)

OpenStudy (zarkon):

this \[f(x)=x-\frac{2}{x^2}\]

OpenStudy (anonymous):

Yes that is what the equation is...

OpenStudy (zarkon):

\[\lim_{x\to\infty}\left(x-\frac{2}{x^2}\right)=\infty\] therefore there is no absolute maximum.

OpenStudy (anonymous):

Thank you!!!